# Any-point Trajectory Modeling for Policy Learning

### 28 Dec 2023 v1

@@ 5.4 Analysis
Figure 5: Success rate of our policy trained with 4%, 10% and 20% action-labeled demos. Our policy trained with only 4% demos performs comparable to BC baseline with 20% demos on LIBERO-Spatial, Object and GOAL, and even better on LIBERO-Spatial. When trained on 20% demos, our performance approaches BC with all training data.
Effect of the number of expert demos. As shown in Figure 5, we evaluated the performance of ATM with various numbers of action labelled demonstrations. In our default setting, as described in Section 5.3, we used 20% of the LIBERO dataset as action-labeled demonstrations. We also reduced the training set further to 10% and 4%, which are equivalent to 5 and 2 demonstrations per task, respectively. The results show a decline in success rates with fewer demonstrations, underscoring the importance of adequate data for training effective low-level control policies. However, ATM’s performance with only 4% training demonstrations is comparable to BC w/ 20% on LIBERO-Object, Goal, and Long, and it is even significantly better on LIBERO-Spatial. Moreover, the performance of ATM trained with 20% of the demonstrations nearly matches that of BC trained with the entire dataset on three short-horizon suites. For long-horizon tasks in LIBERO-Long, enhanced by the fine-grained subgoals provided by predicted tracks, ATM with 20% demonstrations significantly outperforms BC w/ 100%. This indicates that pre-training on an action-free video dataset using our any-point trajectory modeling framework notably improves the sample efficiency of policy learning.
Figure 6: Success rate with different track lengths. In general, longer tracks correspond to higher success rates, except for LIBERO-Object which performs best with 8-step tracks.
@@
Table 1: Ablation study on image masking of track transformer, where “w/o image masking” represents that we do not mask out image patches during track transformer training and “w/ image masking” means we randomly mask 50% patches. We can see that mask image modeling in track transformer improves the policy performance.
@@
Table 2: Ablation study on the policy architecture. We explore the effect of the tracks fed into the policy in two positions: spatial transformer (early fusion) and MLP head (late fusion), as illustrated in Figure 9.
@@
Effect of track length. The track length of our pre-trained track transformer is 16 and we utilize all 16 points of each track by default as input for our policy. In this experiment, we investigate the effect of track length of our policy input, by training our policy with the predicted tracks truncated into 4, 8, and 16 steps. The results in Figure 6 demonstrate that longer tracks result in higher success rates in general, except LIBERO-Object where track length $=8$ yields the best performance. In particular, reducing track length leads to larger performance drops on LIBERO-Goal and LIBERO-Long. These tasks require a comprehensive understanding of the task goal. Longer tracks provide more detailed and specific subgoals for each task, which is crucial for guiding the agent’s movements and actions in subsequent stage. Conversely, for LIBERO-Object which emphasizes precision operations over understanding of the task goal, a policy with a length of 16 underperforms slightly compared to the model with a length of 8. We hypothesize that the longer tracks might interfere with the learning of inverse dynamics due to noise. This also supports our approach of employing tracks both as subgoal conditions and in leveraging the nature of inverse dynamics.
Figure 7: The attention maps of BC and Ours in spatial transformer. We extract the attention weights between spatial CLS token and RGB tokens, highlighting the policy’s focus on specific spatial regions during decision-making. The heatmaps reveal our policy’s targeted attention on task-relevant areas, in contrast to BC’s tendency to focus on irrelevant backgrounds. This underscores the effectiveness of input tracks in the spatial transformer as good task prompts, guiding the CLS token to attend to appropriate areas.
@@
Effect of image masking in track transformer. we investigate the effects of image patch masking within our track transformer on policy performance. We conduct an ablation study by nullifying the image patch masking (setting the masking ratio to 0%) and comparing it against our standard configuration (50% masking ratio). In Table 1, the results reveal a slight decline in policy performance across three out of four test suites (LIBERO-Object, LIBERO-Goal, and LIBERO-Long) when image masking is omitted. This outcome suggests that reconstructing masked image patches is a good auxiliary task for the track transformer.
@@
Effect of early and late fusion in policy architecture. As shown in Figure 3 and Figure 9, the predicted tracks are fed into the policy from two places: spatial transformer (early fusion) and MLP head (late fusion). We conduct ablation studies by removing these two track inputs to illustrate the effect of them respectively. The results are shown in Table 2. We can see that removing the late fusion leads to the most significant performance drop; on LIBERO-Goal, w/ only early fusion performs similarly to other baselines, whereas late fusion performs marginally worse than our full method. This reinforces the strength of tracks of specifying subgoals, and suggests that unencoded tracks themselves are easily interpretable, enabling goal-following behavior.
5.5 Visualization
Figure 8: We visualize the predicted tracks during online rollout in these three example tasks. Each subfigure includes a title of the benchmark’s language description to define the task goal, base-view and wrist-view images with corresponding tracks, and our annotations detailing key events (not model inputs). We can see that predicted tracks align with language instructions throughout the process. It decomposes the high-level task goal into fine-grained subgoals at each timestep. And the policy correctly follows the subgoals, leading to the successful completion of the tasks.
Attention map visualization. To demonstrate how tracks guide the policy, we visualize the attention maps between the spatial CLS token and RGB tokens in spatial transformer of BC and our method. Figure 7 demonstrates that our method effectively focuses on the relevant spatial regions, as specified by the textual instructions. Specifically, it attends to the cream cheese, bowl, and wine bottle in the respective example tasks, while BC is usually distracted by the irrelevant regions, highlighting the superior capability of the tracks as better task prompts.
Tracks prediction visualization. To qualitatively illustrate the effectiveness of our track transformer, we visualized the predicted tracks during an online rollout. As depicted in Figure 8, these tracks consistently align with the provided task-specific language instructions, maintaining directional accuracy throughout. Moreover, the predicted tracks effectively decompose the abstract textual instructions into detailed subgoals. This decomposition, combined with the policy’s precise lane adherence, contributes significantly to its promising performance.
6 Conclusions
@@
In this work, we present an any-point trajectory modeling framework as video pre-training that effectively learns behaviors and dynamics from action-free video datasets. After pre-training, by learning a track-guided policy, we demonstrate significant improvements over prior state-of-the-art approaches for video pre-training followed by policy learning. We show that a particle-based representation is interpretable, structured, and naturally incorporates physical inductive biases such as object permanence. We hope our works will open doors to more exciting directions in learning from the rich source of video data for robotics applications.
7 Limitations
One limitation of our current work is that we only pre-train on in-domain videos, limiting the type of variations our track-guided policy could generalize to. In the future, we hope to scale up our work to more diverse video datasets. Additionally, even after we have pre-trained a track model, currently we are still relying on action-labeled demonstration data to learn the control policies. In the future, we hope to extend our work to reinforcement learning settings where no additional demonstrations are needed in policy learning.
8 Acknowledgement
@@
Appendix
@@
Appendix A Additional Experimental Results
A.1 Numerical Results
[…]
Table 3: Average success rate on LIBERO benchmark. Our method performs consistently better than all the baselines across all suites. UniPi-Replan is only evaluated for a single seed due to the computation cost.
@@ A.1 Numerical Results
Ours $68.50\pm 1.78$ $\bf 68.00\pm 6.18$ $\bf 77.83\pm 0.82$ $\bf 39.33\pm 15.80$ $\bf 48.41\pm 2.09$
A.2 Comparisons to Variations of the UniPi Baseline
In order to decouple the two advantages of ATM over UniPi additionally compare with a UniPi variation where we train the video prediction model to predict a fixed time step into the future $\tilde{o}_{t+H}=f_{\theta}(o_{t})$ for every $H$ steps of policy execution, where $H=8$. As this variation replans the sub-goal more frequently, we call this method UniPi-Replan, similar to the implementation in [4]. The results are shown in Table 3. Surprisingly, we found that this variation performs even worse than UniPi. We thus draw the conclusion that a structured sub-goal representation can be much more effective than an image sub-goal. The reason is that predicting an image goal at a fixed future time step can be a difficult objective for the video prediction model, leading to inconsistent subgoals. Please refer to the failure videos of various baselines on our website. Additionally, this method requires heavy computation. A comparison of the computation needed is shown in Table 4. Due to the computation cost, we only evaluate UniPi-Replan on LIBERO-Object and LIBERO-Goal and report the average success across 10 trials on one policy seed.
Table 4: Computation cost for performing a single task consisting of 600 time steps for different methods. ATM performs trajectory generation at each time step while being computationally efficient. UniPi performs an open-loop future goal generation once at the beginning using a video diffusion model. UniPi-Replan generates image sub-goals at more frequent time steps and is the slowest.
Method ATM UniPi UniPi-Replan
Computation (TFLOPS) 1.56 39.29 2946.75
@@ A.2 Comparisons to Variations of the UniPi Baseline
Appendix B Detailed Policy Architecture
Spatial Encoding: We encode all modalities at each timestep. We first leverage the frozen track transformer to propose a set of tracks as described in Section 4.1 for all $V$ frames in $o_{t}$. We then project each modality with modality-specific encoders into a shared embedding space $\mathbb{R}^{D}$. Each modality’s tokens are embedded using a shared learned modality token, and modality-specific positional embeddings. We then concatenate tokens across all modalities and views with a learned spatial CLS token, and perform self-attention on the sequence. We extract the spatial CLS token as the representation.
Appendix C Implementation and Experimental Details
C.1 Track Filtering for Track Transformer Training
As introduced in Section 4.1, we adopt a heuristic to filter out the static points in the background and then utilize an off-the-shelf tracking model to generate the corresponding tracks of the large-motion points. The visualization of the sampled points before and after the filtering process is shown in Figure 10.
Figure 10: Given a video (left), we query 1000 randomly sampled points using an off-the-shelf TAP model (middle), where each colored dot represents the starting position of a track. We then filter the tracks using a heuristic of their position displacement across the video and re-sample about these points (right). From observation, the extracted tracks are concentrated around informative objects, such as robot’s gripper and manipulation targets.
C.2 Training Details
We list the training hyperparameters for the track transformer and track-guided policy in Table 6 and 6, which are fixed for all experiments in Section 5. We train all models on 4 A100 GPUs with DeepSpeed strategy.
We train the track transformer using the ground truth tracks generated by CoTracker [18] and save the checkpoint with the lowest validation loss as our final model to apply for policy learning. We do not incorporate frame stacking for track transformer to avoid the copycat problem (or named causal confusion) [7, 44, 45].
Table 5: Hyperparameters of track transformer training.
Table 6: Hyperparameters of policy training.
@@ C.2 Training Details
Table 6: Hyperparameters of policy training.

### 16 Feb 2024 v2

@@ V-C Ablation Analysis
Figure 8: Success rate of our policy trained with 4%, 10% and 20% action-labeled demos. Our policy trained with only 4% demos performs comparably to BC baseline with 20% demos on LIBERO-Spatial, Object, and GOAL, and even better on LIBERO-Spatial. When trained on 20% demos, our performance approaches BC with all training data.
[…]
Figure 9: We plot the success rates of the policies learned with predicted trajectories of different lengths. Generally, longer trajectory length improves the performance, but the benefit tends to plateau at 16.
@@
TABLE II: Ablation study on image masking of track transformer, where “w/o image masking” represents that we do not mask out image patches during track transformer training and “w/ image masking” means we randomly mask 50% patches. We can see that mask image modeling in track transformer improves the policy performance.
@@
TABLE III: Ablation study on the policy architecture. We explore the effect of the tracks fed into the policy in two positions: transformer input (early fusion) and MLP head (late fusion), as illustrated in Figure 3.
@@
Effect of track length. The track length of our pre-trained track transformer is 16 and we utilize all 16 points of each track by default as input for our policy. In this experiment, we investigate the effect of track length of our policy input, by training our policy with the predicted tracks truncated into 4, 8, and 16 steps. The results in Figure 9 demonstrate that longer tracks result in higher success rates in general, except LIBERO-Object where track length $=8$ yields the best performance. In particular, reducing track length leads to larger performance drops on LIBERO-Goal and LIBERO-Long. These tasks require a comprehensive understanding of the task goal. Longer tracks provide more detailed and specific subgoals for each task, which is crucial for guiding the agent’s movements and actions in subsequent stage. Conversely, for LIBERO-Object, which emphasizes precision operations over understanding of the task goal, a policy with a length of 16 underperforms slightly compared to the model with a length of 8. We hypothesize that the longer tracks might interfere with the learning of inverse dynamics due to noise. This also supports our approach of employing tracks both as subgoal conditions and in leveraging the nature of inverse dynamics.
@@
Effect of image masking in track transformer. When training the track transformer, we randomly mask out patches in the images and learn to reconstruct them as an auxiliary task. We conduct an ablation study by removing the image masking loss and comparing it against our standard configuration, where we randomly mask out $50\%$ of the image masks. In Table II, the results reveal a slight decline in policy performance when image masking is omitted, suggesting image masking can be a useful auxiliary task.
@@
Effect of early and late fusion in policy architecture. As shown in Figure 3, the predicted tracks are fed into the policy both before and after the transformer architecture within the policy, which we call early fusion and late fusion respectively. We conduct ablation studies by removing these two track inputs. The results are shown in Table III. We can see that removing the late fusion leads to the most significant performance drop; on LIBERO-Goal, w/ only early fusion performs similarly to other baselines, whereas only late fusion performs marginally worse than our full method. This suggests a late fusion of the predicted tracks acts as useful subgoals that help the policy better understand the tasks in a multi-task learning setting. The subgoal prediction is more robust as it is trained on a larger video dataset.
VI Conclusions
@@
In this work, we present an any-point trajectory modeling framework as video pre-training that effectively learns behaviors and dynamics from action-free video datasets. After pre-training, by learning a track-guided policy, we demonstrate significant improvements over prior state-of-the-art approaches and show effective learning from out-of-distribution human videos. We show that a particle-based representation is interpretable, structured, and naturally incorporates physical inductive biases such as object permanence. We hope our works will open doors to more exciting directions in learning from videos with structured representations.
One limitation is that our approach still relies on a set of action-labeled demonstration trajectories for mapping to actions, which limits the generalization of the learned policies. Future works can consider learning the trajectory-following policies using reinforcement learning so that no additional demonstration data are needed. Another limitation of our method is that the video dataset we use in this paper only contains small domain gaps. Learning from in-the-wild video dataset poses additional challenges such as multi-modal distribution, diverse camera motions, and sub-optimal motions. We leave these extensions for future work.
VII Acknowledgement
@@
Appendix A Additional Experimental Results
A-A Simulation Experiments
@@
Attention map visualization. To demonstrate how tracks guide the policy, we visualize the attention maps between the spatial CLS token and RGB tokens in the spatial transformer of BC and our method. Figure 10 demonstrates that our method effectively focuses on the relevant spatial regions, as specified by the textual instructions. Specifically, it attends to the cream cheese, bowl, and wine bottle in the respective example tasks, while BC is usually distracted by the irrelevant regions, highlighting the superior capability of the tracks as better task prompts.
A-B Discussions on the UniPi Baselines
TABLE IV: Computation cost for performing a single task consisting of 600 time steps for different methods. ATM performs trajectory generation at each time step while being computationally efficient. UniPi performs an open-loop future goal generation once at the beginning using a video diffusion model. UniPi-Replan generates image sub-goals at more frequent time steps and is the slowest.
Method ATM UniPi UniPi-Replan
Computation (TFLOPS) 1.56 39.29 2946.75
In order to decouple the two advantages of ATM over UniPi additionally compare with a UniPi variation where we train the video prediction model to predict a fixed time step into the future $\tilde{o}_{t+H}=f_{\theta}(o_{t})$ for every $H$ steps of policy execution, where $H=8$. As this variation replans the sub-goal more frequently, we call this method UniPi-Replan, similar to the implementation in [5]. The results are shown in Table V. Surprisingly, we found that this variation performs even worse than UniPi. We thus draw the conclusion that a structured sub-goal representation can be much more effective than an image sub-goal. The reason is that predicting an image goal at a fixed future time step can be a difficult objective for the video prediction model, leading to inconsistent subgoals. Please refer to the failure videos of various baselines on our website. Additionally, this method requires heavy computation. A comparison of the computation needed is shown in Table IV. Due to the computation cost, we only evaluate UniPi-Replan on LIBERO-Object and LIBERO-Goal and report the average success across 10 trials on one policy training random seed.
Figure 10: The attention maps of BC and Ours in the spatial transformer. We extract the attention weights between spatial CLS tokens and RGB tokens, highlighting the policy’s focus on specific spatial regions during decision-making. The heatmaps reveal our policy’s targeted attention on task-relevant areas, in contrast to BC’s tendency to focus on irrelevant backgrounds. This underscores the effectiveness of input tracks in the spatial transformer as good task prompts, guiding the CLS token to attend to appropriate areas.
TABLE V: Average success rate on LIBERO benchmark. Our method performs consistently better than all the baselines across all suites. UniPi-Replan is only evaluated for a single seed due to the computation cost.
@@ A-B Discussions on the UniPi Baselines
ATM (Ours) $68.50\pm 1.78$ $\bf 68.00\pm 6.18$ $\bf 77.83\pm 0.82$ $\bf 39.33\pm 15.80$ $\bf 48.41\pm 2.09$
@@
Figure 12: Given a video (left), we query 1000 randomly sampled points using an off-the-shelf TAP model (middle), where each colored dot represents the starting position of a track. We then filter the tracks using a heuristic of their position displacement across the video and re-sample around these points (right). We can see that extracted tracks are concentrated around informative objects, such as the robot’s gripper and manipulation targets.
A-C Human-to-robot Transfer Details
To demonstrate the potential of ATM to leverage out-of-domain videos, we design human-to-robot transfer tasks in three different settings: 1) deformable object: fold the cloth and pull it to the right, 2) long horizon: put the tomato into the pan and close the cabinet door, and 3) tool using: use the broom to sweep the toys into the dustpan and put it in front of the dustpan. We collect 10 robot teleportation trajectories (action-labeled) and 100 human manipulation videos (action-free). We compare three methods: 1) behavioral cloning only with 10 action-labeled robot demos, 2) ATM, training track transformer with only 10 robot demos, and 3) ATM, training a track transformer with both action-free human and action-labeled robot data. For each task, we train each method with three different random seeds, evaluate them in the real world 10 times, and report the average success rate in Table I.
Human-to-Robot visualization. The detailed video visualizations of human-to-robot skill transfer are shown in Figure 13. Due to limited training samples, the track transformer trained with only 10 robot videos fails to predict the future point trajectories, leading to low success rates. In contrast, the trajectory model trained with human videos generates high-quality tracks in each frame, guiding the agent to successfully complete the task with only 10 action-labeled trajectories. This indicates our any-point trajectory modeling enables to transfer the motion prior from cross-embodiment videos to robot skills, which significantly improves policy learning.
TABLE VI: Hyperparameters of track transformer training.
TABLE VII: Hyperparameters of policy training.
@@ A-C Human-to-robot Transfer Details
TABLE VII: Hyperparameters of policy training.
Figure 13: The visualizations of human demos and rollout videos of ATM policies trained with and without human data. We can see that ATM is able to take advantage of out-of-domain videos, i.e., human videos, to generate more precise tracks, resulting in better policy performance.
Appendix B Implementation Details
B-A Policy Architecture
Spatial Encoding: We encode all modalities at each timestep. We first leverage the frozen track transformer to propose a set of tracks for all $V$ frames in $o_{t}$. We then project each modality with modality-specific encoders into a shared embedding space $\mathbb{R}^{D}$. Each modality’s tokens are embedded using a shared learned modality token, and modality-specific positional embeddings. We then concatenate tokens across all modalities and views with a learned spatial CLS token, and perform self-attention on the sequence. We extract the spatial CLS token as the representation.
B-B Efficient Training with Point Filtering
We adopt a heuristic to filter out the static points in the background and then utilize an off-the-shelf tracking model to generate the corresponding tracks of the large-motion points. The visualization of the sampled points before and after the filtering process is shown in Figure 12.
B-C Training Details
We list the training hyperparameters for the track transformer and track-guided policy in Table VII and VII, which are fixed for all experiments on LIBERO benchmark. We train all models on 4 A100 GPUs with DeepSpeed strategy.
We train the track transformer using the ground truth tracks generated by CoTracker [18] and save the checkpoint with the lowest validation loss as our final model to apply for policy learning. We do not incorporate frame stacking for track transformer to avoid causal confusion [48].